An undivided multilane highway segment has two 11-ft lanes in the eastbound direction with no shoulders and a 55mph speed limit. This highway segment has 27 access points on a 0.625 mile,4% upgrade. The peak hour factor is 0.9,the directional peak hour volume is 2880 vehicles/hr which includes 200 single unit trucks (SUTs) and 60 tractor trailers (TTs). What are the estimated speed, density, and level of service of this upgrade?
Answer #1

Answer:

Given that

Base free flow speed (BFFS) = 55 mph

Lane width = 11 ft

Reduction in speed corresponding to lane width, fLW = 1.9 mph

Access points = 17 / 0.625 miles = 27.2 /mile

Reduction in speed corresponding to access points, fID = 5.0 mph

No. of lanes = 2

Reduction in speed corresponding to undivided highway, fN = 1.6 mph

Free Flow Speed (FFS) = BFFS – fLW – fN – fID = 55 – 1.9 – 1.6 – 5 = 46.5 mph

Peak Flow, V = 2880 veh/hr

Peak-hour factor = 0.9

SUT = 200 = 6.9 %

TT = 60 = 2.1%

4% upgrade for a distance of 0.625 miles

fHV = 1/ (1 + 0.069 (2-1) + 0.021 (3-1)) = 1/1.111 = 0.90

fP = 1.0

Peak Flow Rate, Vp = V / (PHV*n*fHV*fP) = 2880/ (0.90*2*0.90*1.0) = 1777.77 ~ 1778 veh/hr/ln

Estimated Speed, S = FFS = 46.5 mph

Density = Vp/S = (1778) / (46.5) = 38.24 veh/mi/ln

Hence Level of Service is E

Density of LOS E should lie between 35 – 43 veh/mi/ln